Nitrogen is throttled at a rate of 0.9 kg/s from a pressure of 10 bar to 1.5 bar. 

Outlet diameter is 10 cm. If outlet temperature was 37oC, calculate velocity at 

outlet and deduce temperature at inlet. Find work and heat exchanged per 

second during this process.


1
Expert's answer
2021-07-05T16:08:05-0400

Gives


"\frac{dm}{dt}=0.9kg/sec,P_2=10bar,P_2=1.5bar,r=5cm,T_1=37°c"

We know that

"\frac{dQ}{dt}=Av"

"v=\frac{\frac{dQ}{dt}}{A}"

Put value


"A=\pi r^2=3.14\times0.05^2=7.85\times10^{-3}m^2"

"v=\frac{0.9}{7.85\times10^{-3}}=114.64m/sec"